OpenStudy (anonymous):

The number of blueberry muffins that a baker makes each day is 40% of the total number of muffins she makes. On Monday, the baker makes 36 blueberry muffins. What is the total number of muffins that the baker makes on Monday? A. 14 B. 9 C. 1440 D. 90

OpenStudy (anonymous):

$36\left(40\right)$36(40) $\mathrm{Remove\:parentheses}:\quad\left(a\right)=a$Remove parentheses: (a)=a $=36\cdot\:40$=36· 40 $\mathrm{Multiply\:the\:numbers:}\:36\cdot\:40=1440$Multiply the numbers: 36· 40=1440 $=1440$=1440

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

$=36\cdot\:0.4$=36· 0.4 $\mathrm{Multiply\:the\:numbers:}\:36\cdot\:0.4=14.4$Multiply the numbers: 36· 0.4=14.4 $=14.4$=14.4

OpenStudy (anonymous):

here is the correct 1